For the circuit shown, find the impedance and phas
Answer #1

Inductive reactance=X_L=\omega L=2\pi f L=2\pi \times 240\times 25\times 10^{-3}=37.7\Omega

Capacitive reactance =X_C=\frac{1}{\omega C}=\frac{1}{2\pi f C}=\frac{1}{2\pi \times 240\times 15\times 10^{-6}}=44.21\Omega

So impedance of LCR circuit is

Z=\sqrt{R^2+(X_L-X_C)^2}=\sqrt{15^2+(37.7-44.21)^2}=16.35\Omega

Phase is

cos\phi=\frac{R}{Z}\\ \implies \phi=cos^{-1}\left(\frac{R}{Z} \right )\\ \implies \phi=cos^{-1}\left(\frac{15}{16.35 } \right )\\ \implies \phi=23.44\degree

For Resistor

I_{rms}=\frac{V}{Z}=\frac{20}{16.35}=1.22A

V_{rms}=I_{rms}R=1.22\times 15=18.3V

P=VI=18.3\times1.22=22.326W

For inductor

I_{rms}=\frac{V}{Z}=\frac{20}{16.35}=1.22A

V_{rms}=I_{rms}X_L=1.22\times 37.7=45.994V

P=VI=45.994\times1.22=56.11W

For Capacitor

I_{rms}=\frac{V}{Z}=\frac{20}{16.35}=1.22A

V_{rms}=I_{rms}X_C=1.22\times 44.21=53.94V

P=VI=53.94\times1.22=65.61W

Phasor diagram

V R

Circuit is capactive